A community in Nevada has 9751 households, with a median house price of $320,000 and a mean price of $392,059. a. Why is the mean larger than the median? The mean is larger than the median because house values tend to be b. The property tax rate is about 1.15%. What total amount of taxes will be assessed on these houses? Round your answer to the nearest integer. The total property taxes is dollars. c. What is the average amount of taxes per house? Round your answer to 2 decimal places. The average amount of taxes per house is Show Work is REQUIRED for this question: Qpen Show Work dollars per house.
Answer #1

a.

We~know,~mean>median~implies~the~distribution~is~positively~skewed~\\~so~the~longer~tail~of~the~distribution~is~towards~higher.~Hence\\ the~mean~is~larger~than~the~median~because~house~values~tend~to~be~larger.

b.

b.~Total~taxes=1.15*9751*392059/100=43964124~dollars\\ c.~Average~amount~of~taxes~per~house=Total~Taxes/total~no.~of~household\\ =43964124/9751=4508.68~dollars